OpenStudy (anonymous):

Is there any way to integrate the following: The indefinite integral of: square root of the fraction: [4612 - 32y]/[4356 - 32y] dy

OpenStudy (anonymous):

Wow...this would take forever. Just have a look here. http://www.wolframalpha.com/input/?i=integrate+square+root+of+%5B4612+-+32y%5D%2F%5B4356+-+32y%5D You can press on "show steps".

OpenStudy (anonymous):

that integral is of the form: \[\large \int\limits \frac {a - by}{c-by}dy= \int\limits \frac {a}{c-by}dy-\int\limits \frac{by}{c-by}dy\]

OpenStudy (anonymous):

first one is just ln, second you can divide... it just big a, b, c's isn't it? or am i way off?